# Membership Privacy Risks of Sharpness Aware Minimization

### v3

@@ Abstract
Optimization algorithms that seek flatter minima such as Sharpness-Aware Minimization (SAM) are widely credited with improved generalization. We ask whether such gains impact membership privacy. Surprisingly, we find that SAM is more prone to membership inference attacks than classical SGD across multiple datasets and attack methods, despite achieving lower test error. This is an intriguing phenomenon as conventional belief posits that higher membership privacy risk is associated with poor generalization. We conjecture that SAM is capable of memorizing atypical subpatterns more, leading to better generalization but higher privacy risk. We empirically validate our hypothesis by running extensive analysis on memorization and influence scores. Finally, we theoretically show how a model that captures minority subclass features more can effectively generalize better and have higher membership privacy risk.
@@ 1 Introduction
Sharpness-Aware Minimization (SAM) has emerged as a prominent optimization technique for improving generalization in deep learning by encouraging flatter minima – i.e., similar loss values for weight perturbations of certain degree around the optima – in the loss landscape (Norton & Royset, 2021; Foret et al., 2020; Wu et al., 2020; Kim et al., 2022; Du et al., 2022; Kwon et al., 2021). Flatter optima are associated with robustness to weight perturbations and have been linked to reduced overfitting and improved test performance (Chen et al., 2023; Foret et al., 2020; Baek et al., 2024).
Models that generalize well are thought to rely less on memorizing specific training examples, which should also improve privacy. Consider membership inference attacks (MIAs) in which an attacker exploits the model confidence gap between training and unseen data to infer if a data point was part of the training data or not (Shokri et al., 2017). Intuitively, when a model strongly overfits (training error $\ll$ test error), MIA would become easier. Yeom et al. (2018) formally showed that, under certain assumptions, the advantage of a threshold-based MIA is upper bounded by the model’s generalization error. In light of this conventional wisdom, one would naturally expect that a technique like SAM – which demonstrably improves generalization – should also decrease a model’s susceptibility to MIAs.
Our work uncovers a striking contradiction to this expectation: we find that models trained with SAM are actually more vulnerable to MIAs than SGD consistently across diverse datasets and attack methods, even as they achieve better generalization (see Tables˜1 and 3). This result prompts a deeper investigation into the relationship between generalization, memorization, and privacy to unearth this phenomenon both empirically and theoretically.
Table 1: Attack accuracy of different types of MIA on SGD, SAM and Sharp showing tradeoffs in test accuracy and MIA attacks. SAM is more prone to MIA attack. In blue we highlight the best performing model on the test set and in red the model against which MIA is more successful.
@@
MIA fundamentally relies on the idea that models behave differently on data points that were part of the training set versus those that were not. This gap is often attributed to overfitting, but it can be more precisely quantified using memorization scores (Feldman, 2020; Feldman & Zhang, 2020), defined via Leave-One-Out (LOO) error. Memorization scores measure the change in model performance when a specific training sample is removed, and thus serve as a proxy for how much the model has memorized that sample. Motivated by this connection, we analyze the memorization scores of SAM-trained models and find that SAM exhibits higher memorization scores than SGD, indicating stronger reliance on individual training samples. While this increased memorization provides a plausible explanation for SAM’s heightened vulnerability to MIA, it raises a key question: "How can a model that memorizes more generalize better?"
@@
We hypothesize that the answer lies in what is being memorized. Under overparameterization (Allen-Zhu et al., 2019), models can memorize not only noise in the data, but also atypical patterns in under-represented subpopulations—i.e., few white tiger images with numerous yellow tiger images. This distinction is important as real world datasets are known to have a long tail of such rare subclasses (Feldman, 2020). We conjecture that SAM is capable of doing more structured memorization, selectively focusing on atypical subclass patterns, which contributes positively to generalization. Corroborating this hypothesis, we observe that SAM’s memorization score distribution is concentrated in the mid range, rather than the high end—which is typically associated with noise memorization (refer to Section˜4.1). This suggests that SAM emphasizes samples that are neither trivially learned nor purely noisy, but instead represent rare, but generalizable sub-patterns.
@@
To further validate this finding, we analyze influence scores—which measure the impact of individual training samples on test predictions (see Section˜4.2). Our results show that, for SAM, samples corresponding to moderate memorization exert higher influence on test predictions compared to SGD, confirming that SAM’s generalization gains derive from its ability to better memorize rare sub-patterns rather than noise. This property benefits generalization, but comes at the cost of increased membership privacy risk due to the increased retention of information an attacker can exploit.
@@
We support our intuition further by introducing a novel metric that quantifies the degree of memorization involved in predicting a test sample in Section˜4.3. Using this metric, we dissect SAM’s performance gains and identify that SAM’s improvements mostly stem from its performance on atypical test samples that depend heavily on a handful of memorized training points. Meanwhile, SGD performs better on typical samples that rely more on broadly learned features, once again confirming the increased MIA risks associated with SAM.
Finally, we complement our empirical findings with a theoretical analysis (Section˜5) that formalizes the relationship between generalization and MIA risk. We consider a data setup comprising (i) majority subclass, (ii) minority subclass, and (iii) pure noise samples, and show that even a linear classifier model that better captures minority subclass features can achieve higher generalization and higher vulnerability to MIA. This theoretical framework provides insight into when and why the conventional belief that better generalization implies lower privacy risk can break down.
@@
In summary, our contributions are the following: (i) we are the first to empirically demonstrate that SAM-trained models exhibit higher membership privacy risk than SGD-trained models, challenging the conventional assumption that improved generalization implies stronger privacy; (ii) we compute and release memorization and influence scores for SAM-trained models; (iii) we offer a detailed and conceptually grounded analysis of the root causes of SAM’s generalization-memorization relationship; (iv) we introduce a novel methodology to dissect generalization gains, proving that SAM’s generalization gains stem from its performance on unseen atypical samples; and (v) we present a theoretical formulation showing that, under a mixture data distribution, a model can simultaneously achieve high generalization and high MIA.
@@ 2.1 Membership Inference attacks
Let’s say that a classifier $f_{v}(\mathbf{x};\theta)$ was trained using dataset $D\sim\mathcal{D}$. We call this learnt model as the victim model. There is an attacker who is given a sample data point $(\mathbf{x},y)$ and infers whether $(\mathbf{x},y)$ $\in$ $D$ (member) or $(\mathbf{x},y)$ $\notin$ $D$ (non-member). The attacker trains a model $f_{a}(\mathbf{x},y)\in\{0,1\}$ that outputs 1 for member and 0 for non-member. Based on the attacker’s knowledge about the victim model, there are many variations of the attack (Hu et al., 2022). In this paper, we consider a black-box attack scenario assuming the attacker only has access to the output probabilities of the victim model.
The most commonly used MIA is directly querying the target sample and using the statistics returned by the model to predict members and non-members of the training data with reasonable accuracy (Shokri et al., 2021; Murakonda & Shokri, 2020; Nasr et al., 2019; Zhang et al., 2021; Long et al., 2017; Sablayrolles et al., 2019; Yeom et al., 2020). Further details are provided in Section˜D.2.
Given $n_{\mathrm{m}}$ member data points and $n_{\mathrm{nm}}$ non-member data points, the empirical accuracy of a MIA is defined as the average of (i) the fraction of member samples correctly predicted as members, and (ii) the fraction of non-member samples correctly predicted as non-members. Formally:
$\mathrm{Acc}_{\mathrm{MIA}}=\frac{1}{2n_{\mathrm{m}}}\sum_{i=1}^{n_{\mathrm{m}}}\big[f_{\mathrm{a}}(\mathbf{x}_{i},y_{i})=1\;\mid\;(\mathbf{x}_{i},y_{i})\in D\big]+\frac{1}{2n_{\mathrm{nm}}}\sum_{j=1}^{n_{\mathrm{nm}}}\big[f_{\mathrm{a}}(\mathbf{x}_{j},y_{j})=0\;\mid\;(\mathbf{x}_{j},y_{j})\notin D\big].$ (4)
@@ 3 Privacy Risks of SAM
Inspired by the link between SAM and generalization and how MIAs should exploit poor generalization, we here scrutinize the membership privacy risk of SAM by comparing the membership attack accuracy (see Equation˜4) of different MIAs against SAM- and SGD-trained models across five different benchmark datasets. Moreover, to further investigate the connection between flatness of minima and membership privacy, we also design a novel optimizer, namely Sharp, that explicitly aims to find sharper minima. The objective function of Sharp is,
This objective can be seen as minimizing the loss at current $w$ while maximizing the loss in the vicinity. We empirically verify that this objective does lead to a sharper minima measuring its hessian trace. Results and discussion about Sharp are available in Section˜F.4.
@@
We use CIFAR-10, CIFAR-100 and Purchase-100 along with two medical datasets Texas-100 and EyePacs. We follow Tang et al. (2022) to determine the partition between training and test data and to determine the subset that constitutes the attacker’s prior knowledge  [^1] We adopt and extend the code in https://github.com/inspire-group/MIAdefenseSELENA. More details about the datasets and the experimental setup can be found in Appendix˜E.
@@ Methods
For each setting, we train a set of models and choose the one achieving highest validation accuracy. We then employ different MIA methods – namely NN-based, confidence-based, entropy-based and modified entropy-based attacks (see Section˜D.2 for a detailed formulation of each MIA) – to evaluate the attack accuracy on the target model. All details about the experimental settings can be found in Appendix˜F.
@@ Results
We report the attack accuracy and test accuracy for each model in Table˜1. We report the mean accuracy over 5 randomized runs with different attack data splits. We observe that while SAM achieves the highest generalization performance, it also incurs in the highest attack accuracy for most datasets and attack methods. Moreover, the MIA gap is more pronounced for image datasets like CIFAR-10, CIFAR-100, and EyePacs than for tabular datasets. This finding is consistent with Feldman & Zhang (2020), since image datasets tend to be characterized by more long-tailed distributions composed of multiple atypical subclasses (see Section˜4.3). The results also show that membership inference attack accuracy decreases for models trained with Sharp (Equation˜5) w.r.t. both SAM and SGD. On the other hand, the test accuracy for Sharp is also lower. This result suggests a strong connection between membership privacy and flatter minima in general.
Finally, to confirm that these results are not model-dependent, we report an ablation study in Appendix˜G and verify that similar findings can be observed for different model architectures over the same datasets.
@@ 4 SAM learns Atypical Subclass Features More
To investigate the source of SAM’s increased membership privacy risk, we analyze its optimization behavior through the lens of memorization and sample influence. We follow the procedure of Feldman & Zhang (2020) to compute the memorization and influence scores for SAM-trained models on CIFAR-100. We then compare these scores against the publicly available scores for SGD-trained models on the same dataset [^2] https://pluskid.github.io/influence-memorization/, enabling a direct comparison of sample-level behavior between the two optimizers.
@@ 4.1 SAM Memorizes Atypical Sub-patterns More
A regression analysis over all samples – shown via the green curve – reveals a consistent deviation from the identity – red – line, skewed towards the top-left quadrant. This indicates a systematic increase in memorization for a large subset of samples under SAM. Crucially, this deviation is not concentrated at the high end of the memorization spectrum. This finding – together with the kernel density plot – supports our hypothesis that SAM does not simply memorize noise, but rather focuses on non-dominant, atypical subclass patterns that are underrepresented in the training distribution. Indeed, if SAM were memorizing sample-specific noise, we would expect a sharp concentration of kernel density at the highest end of the spectrum and a deviation of the regression curve in the top-right quadrant.

### v4

@@ Abstract
Optimization algorithms that seek flatter minima, such as Sharpness-Aware Minimization (SAM), are credited with improved generalization and robustness to noise. We ask whether such gains impact membership privacy. Surprisingly, we find that SAM is more prone to Membership Inference Attacks (MIA) than classical SGD across multiple datasets and attack methods, despite achieving lower test error. This suggests that the geometric mechanism of SAM that improves generalization simultaneously exacerbates membership leakage. We investigate this phenomenon through extensive analysis of memorization and influence scores. Our results reveal that SAM is more capable of capturing atypical subpatterns, leading to higher memorization scores of samples. Conversely, SGD depends more heavily on majority features, exhibiting worse generalization on atypical subgroups and lower memorization. Crucially, this characteristic of SAM can be linked to lower variance in the prediction confidence of unseen samples, thereby amplifying membership signals. Finally, we model SAM under a perfectly interpolating linear regime and theoretically show that sharpness regularization inherently reduces variance, guaranteeing a higher MIA advantage for confidence and likelihood ratio attacks.
@@ 1 Introduction
Sharpness-Aware Minimization (SAM) has emerged as a prominent optimization technique for improving generalization in deep learning by encouraging flatter minima – i.e., similar loss values for weight perturbations of certain degree around the optima – in the loss landscape (Norton & Royset, 2021; Foret et al., 2020; Wu et al., 2020; Kim et al., 2022; Du et al., 2022; Kwon et al., 2021). Flatter optima have been linked to robustness to noise and improved test performance (Chen et al., 2023; Foret et al., 2020; Baek et al., 2024), while a tension exists between whether SAM’s implicit bias is geared more towards diversity (Springer et al., 2024) or simplicity (Andriushchenko et al., 2023; Chang & Khanna, 2025) of features.
Models that generalize well are thought to rely less on memorizing specific training examples, which should also improve privacy. Consider membership inference attacks (MIAs) in which an attacker exploits the model behavior gap between training and unseen data to infer if a data point was part of the training data or not (Shokri et al., 2017). Intuitively, when a model strongly overfits (training error $\ll$ test error), MIA would become easier. Yeom et al. (2018) formally showed that, under certain assumptions, the advantage of a threshold-based MIA is upper bounded by the model’s generalization error. In light of this, one would naturally expect that a technique like SAM – which demonstrably improves generalization – should also decrease a model’s susceptibility to MIAs.
Contrary to this expectation, we find that models trained with SAM are actually more vulnerable to MIAs than SGD consistently across diverse datasets and attack methods, even as they achieve better generalization (see Tables˜1 and 2). Furthermore, this finding challenges the notion that flatter minima=good from a privacy standpoint, calling for a deeper investigation into the relationship between generalization, memorization, and privacy to unearth this phenomenon both empirically and theoretically. Our work is the first to systematically demonstrate higher membership privacy leakage for a sharpness-based algorithm known to generalize better, connecting loss sharpness to MIA vulnerability. We note that there have been previous works that exhibit utility–privacy tradeoffs (Long et al., 2018; Carlini et al., 2021; Chen et al., 2022; Liu et al., 2024) or demonstrate that higher generalization does not necessarily decrease privacy leakage (Kaya & Dumitras, 2021; Del Grosso et al., 2023).
Table 1: Attack accuracy of direct threshold MIA on SGD and SAM showing tradeoffs in test accuracy and MIA attacks. In blue we highlight the best performing model on the test set and in red the model against which MIA is more successful. SAM is more prone to direct threshold attacks.
@@
The difference in model behavior on data points that were part of the training set versus those that were not can be more precisely quantified using memorization scores (Feldman, 2020; Feldman & Zhang, 2020), defined via Leave-One-Out (LOO) error. Memorization scores measure the change in model performance when a specific training sample is removed, and thus serve as a proxy for how much the model has memorized that sample. Motivated by this connection, we analyze the memorization scores of samples trained with SAM and find that SAM exhibits more memorization than SGD, indicating a stronger reliance on individual training samples. While this increased memorization provides a plausible explanation for SAM’s heightened vulnerability to MIA, it raises a key question: "How can a model that memorizes more generalize better?"
@@
We hypothesize that the answer lies in what is being memorized. Under overparameterization (Allen-Zhu et al., 2019), models can learn not only noise in the data, but also atypical patterns in under-represented subpopulations through memorization—i.e., few white tiger images with numerous yellow tiger images. This distinction is important as real world datasets are known to have a long tail of such rare subclasses (Feldman, 2020). We conjecture that SAM is capable of doing more structured memorization, selectively focusing on atypical subclass patterns, which contributes positively to generalization. Corroborating this hypothesis, we observe that SAM’s memorization score distribution is concentrated in the mid range, rather than the high end—which is typically associated with noise memorization (refer to Section˜4.1). This suggests that SAM emphasizes samples that are neither trivially learned nor purely noisy, but instead represent rare, but generalizable sub-patterns.
@@
To further validate this finding, we analyze influence scores—which measure the impact of individual training samples on test predictions (see Section˜4.2). Our results show that, for SAM, samples corresponding to moderate memorization exert higher influence on test predictions compared to SGD, confirming that SAM’s generalization gains derive from its ability to better capture rare sub-patterns. Conversely, for SGD, the lower influence of such points implies that the majority pattern is learned dominantly: since this feature is redundant across many samples, the marginal influence of any single point is diluted. These results seem to suggest more that SAM’s implicit bias is towards diversity as opposed to simplicity.
@@
We support our intuition further by introducing a novel metric that quantifies the degree of memorization involved in predicting a test sample in Section˜4.3. Using this metric, we dissect SAM’s performance gains and identify that SAM’s improvements mostly stem from its performance on atypical test samples that depend heavily on a handful of memorized training points. Meanwhile, SGD performs slightly better on typical samples that rely more on broadly learned features. Stronger influence of minority samples can lead to greater membership privacy risk due to the increased retention of information an attacker can exploit.
However, what intrinsic mechanism drives this structured memorization? Results for confidence threshold attack indicates that SGD produces more predictions with extreme confidence compared to SAM. These samples reside beyond the threshold and are source of attacker’s error. This suggests a distinct geometric effect: SAM induces a shrinkage in the variance of the model’s output predictions. This property translates to structured memorization and suppression of majority sub-feature. Relying heavily on a single/few majority subclass feature to classify diverse inputs requires the model to assign large weights to that feature. Geometrically, this creates a steep decision boundary and, consequently, high output variance under perturbation. By penalizing this sharpness, SAM prohibits the amplification of the majority feature, effectively forcing the model to distribute its reliance across diverse, subclass-specific features.
Completing this conjecture, we provide a theoretical foundation for this mechanism in Section˜5. We prove that the interpolating solution favored by sharpness-aware geometry inherently reduces the variance of the output logits. We then demonstrate how this variance reduction amplifies the attacker’s advantage for both confidence-based and likelihood ratio attacks. Our proofs highlight a strong result: SAM is more vulnerable at any fixed threshold. We empirically corroborate this with ROC curves in Figure 7. Lastly, we theoretically analyze a dataset with majority and minority subclasses and show how capturing minority feature better leads to enhanced generalization in Section B.
@@
In summary, our contributions are the following: (i) we are the first to empirically demonstrate that SAM-trained models exhibit higher membership privacy risk than SGD-trained models, serving as a cautionary tale against flatter minima=good notion from a privacy standpoint; (ii) we offer a detailed and conceptually grounded analysis of the root causes of SAM’s generalization-memorization relationship, suggesting SAM’s implicit bias towards diversity; (iii) we introduce a novel methodology to dissect generalization gains, proving that SAM’s generalization gains stem from its performance on unseen atypical samples; (iv) we theoretically show variance shrinkage effect of interpolating sharpness-aware solutions and how it increases MIA risk for both confidence and likelihood ratio attacks; and (v) we theoretically formulate a data distribution composed of subclasses where stronger alignment with minority subclass features enhances generalization.
Table 2: Comparison of online shadow-model MIA on SGD and SAM. In blue we highlight the best performing model on the test set, and in red the model with higher privacy leakage (higher AUC, Attack Accuracy, and TPR@0.1%FPR). SAM is more prone to shadow-model attacks.
SGD SAM
Dataset Attack Test Acc AUC Attack Acc TPR@.1 Test Acc AUC Attack Acc TPR@.1
CIFAR-100 RMIA 67.7% 90.4% 80.8% 21.0% 69.1% 91.6% 82.2% 23.4%
LiRA 92.6% 82.9% 27.0% 93.7% 84.1% 31.0%
CIFAR-10 RMIA 92.3% 71.4% 63.5% 4.8% 93.1% 74.9% 65.9% 6.7%
LiRA 73.0% 64.2% 8.8% 76.4% 66.7% 12.5%
Purchase100 RMIA 76.5% 68.8% 62.7% 1.5% 77.4% 70.2% 63.7% 1.7%
LiRA 68.9% 62.6% 1.4% 70.2% 63.4% 1.6%
Texas100 RMIA 46.9% 79.8% 70.8% 2.9% 49.2% 80.6% 71.5% 2.8%
LiRA 80.8% 71.3% 6.9% 81.6% 72.0% 8.2%
@@ 2.1 Membership Inference Attacks
Consider a victim model $f_{v}$ trained on dataset $D\sim\mathcal{D}$ and attack model $f_{a}$. In a black-box setting, an attacker infers whether a sample $(\mathbf{x},y)$ belongs to $D$ (IN) or not (OUT). In this paper, we consider two types of attacks: Direct threshold attacks, which directly learns a threshold from obtained member/non-member data, such as confidence and entropy attacks; and Shadow model attacks, which train proxy models to calibrate membership scores, such as Likelihood Ratio Attack (LiRA) and Robust Membership Inference Attack (RMIA) (Carlini et al., 2022; Zarifzadeh et al., 2024).
We quantify privacy risk using the empirical attack accuracy, defined as the average of the true positive (TPR) and true negative rates (1-FPR):
$\mathrm{Acc}_{\mathrm{MIA}}=\frac{1}{2n_{\mathrm{m}}}\sum_{i=1}^{n_{\mathrm{m}}}\mathds{1}\big[f_{\mathrm{a}}(\mathbf{x}_{i},y_{i})=1\big]+\frac{1}{2n_{\mathrm{nm}}}\sum_{j=1}^{n_{\mathrm{nm}}}\mathds{1}\big[f_{\mathrm{a}}(\mathbf{x}_{j},y_{j})=0\big],$ (4)
where $n_{m},n_{nm}$ are the counts of IN and OUT samples. Additionally, metrics such as Area Under ROC Curve (AUC) and TPR at low FPR are employed to characterize vulnerability. Further details are provided in Section˜D.2.
@@ 3 Privacy Risks of SAM
Inspired by the link between SAM and generalization and how MIAs should exploit poor generalization, we here scrutinize the membership privacy risk of SAM by comparing the membership attack accuracy (see Equation˜4 and Table˜2) of different MIAs against SAM- and SGD-trained models across five different benchmark datasets for direct threshold attacks and four different benchmark datasets for shadow model attacks.
@@
We use CIFAR-10, CIFAR-100 and Purchase-100 along with two medical datasets Texas-100 and EyePacs. For direct threshold attacks, we follow Tang et al. (2022) to determine the partition between training and test data and to determine the subset that constitutes the attacker’s prior knowledge  [^1] We adopt and extend the code in https://github.com/inspire-group/MIAdefenseSELENA. For shadow model attacks, we use a different dataset split to account for shadow model training  [^2] We adopt and extend the code in https://github.com/orientino/lira-pytorch. More details about the datasets and the experimental setup can be found in Appendix˜G.
@@ Methods
For direct threshold attacks, we train a set of models and choose the one achieving highest validation accuracy. We then employ different MIA methods – namely NN-based, confidence-based, entropy-based and modified entropy-based attacks (see Section˜D.2 for a detailed formulation of each MIA) – to evaluate the attack accuracy on the target model. For shadow model attacks, we generate 256 random half-splits of the training dataset into member set and non-member set and train a model for each split. One model is chosen as the target model and all the other models are used as shadow models for reference. More details about the experimental settings can be found in Appendix˜H.
@@ Results
Direct Threshold Attacks We report the attack accuracy and test accuracy for each model in Table˜1. We observe that while SAM achieves higher generalization performance, it also incurs highest attack accuracy for almost all settings. To further investigate the connection between flatness of minima and membership privacy, we report the results for other sharpness-aware optimizers and custom designed optimizer that explicitly aims to find sharper minima in Section E. The results support a relationship between loss landscape geometry and membership privacy, with other optimizers exhibiting similar behavior. For confidence threshold attack, we observe that SGD model incur higher number of extremely confident predictions compared to SAM for both correctly classified and wrongly classified non-members. Because the threshold is typically set near a high value (i.e. 0.92), non-member samples with high confidence are missed by the attacker. From this observation, we conjecture that variance of the model output is an important factor in MIA risk. To confirm that these results are not model-dependent, we report an ablation study in Appendix˜I and verify that similar findings can be observed for different model architectures over the same datasets.
Shadow Model Attacks Table 2 summarizes the results for online Robust MIA (RMIA) and Likelihood Ratio Attack (LiRA), reporting Attack Accuracy, AUC, and TPR at 0.1% FPR averaged over 10 attack splits (Zarifzadeh et al., 2024; Carlini et al., 2022). Consistent with the previous results, SAM achieves higher generalization while incurring non-trivially higher privacy leakage across nearly all settings. We note that the reported test accuracies differ from Table 1 due to difference in data splits for shadow model training.
The results for offline attacks are presented in Table 4. Analogous to the online setting, SAM exhibits superior generalization paired with heightened MIA vulnerability. Since shadow model attacks represent the state-of-the-art in membership inference, these results suggest that SAM’s vulnerability is not merely an artifact of global threshold shifts. Instead, it points to an intrinsic geometric property of SAM that persists even under rigorous, sample-specific SOTA attacks.
@@ 4 SAM learns Atypical Subclass Features More
To investigate the source of SAM’s increased membership privacy risk, we analyze its optimization behavior through the lens of sample memorization and influence. We follow the procedure of Feldman & Zhang (2020) to compute the memorization and influence scores for SAM-trained models on CIFAR-100. We then compare these scores against the publicly available scores for SGD-trained models on the same dataset [^3] https://pluskid.github.io/influence-memorization/, enabling a direct comparison of sample-level behavior between the two optimizers.
@@ 4.1 SAM Memorizes Atypical Sub-patterns More
A regression analysis over all samples – shown via the green curve – reveals a consistent deviation from the identity – red – line, skewed towards the top-left quadrant. This indicates a systematic increase in memorization for a large subset of samples under SAM. Crucially, this deviation is not concentrated at the high end of the memorization spectrum. This finding – together with the kernel density plot – supports our hypothesis that SAM does not simply memorize pure noise samples, but rather focuses on non-dominant, atypical subclass samples that are underrepresented in the training distribution. Indeed, if SAM were picking up sample-specific noise, we would expect a sharp concentration of kernel density at the highest end of the spectrum and a deviation of the regression curve in the top-right quadrant.